When the amount of (p) is invested at annual interest rate (r) compounded for (t) years, it will grow to. An amount (a) given by the formula
A=P (1+r)^t
1 suppose php 20,000 is invested at annual interest rate(r) compounded annually
In 3 years, it grows to php 23,152. What is the interest rate?
2. Suppose in 10 years. The php 20,000 invested grows to php 35816.95, what is the interest rate?
Hello :)
Hope you can help me solve this problem :)
Im really thankful if anyone can answer it :)

Answer by Boreal(15235)

23152=20000(1+r)^3
divide by 20000
1.1576=(1+r)^3
ln of both sides, remembering that ln of an exponent puts the exponent in front.
0.1463=3 ln (1+r)
divide by 3
0.04878=ln(1+r)
raise both to e power
1.0499=1+r
r=0.0499 or probably 5%.
Check: 20000(1.05^3)=23152
===================================
35816.95=20000(1+r)^10
Do the same way
1.79084=(1+r)^10. Without rounding, take the ln of the left and then divide by 10, since the right will become 10 ln (1+r)
0.0583=ln(1+r)
now raise both to e power
1.05999=1+r
r=6%
20000(1.06)^10=35816.95